# Investigating and Defending Shortcut Learning in Personalized Diffusion Models

### v3

@@ 5.2 Effectiveness, Efficiency, and Faithfulness
Effectiveness Comparison. We present the effectiveness of different purification across six perturbation methods in Table 1. From the table, we can see that compared to the clean case, training on perturbing data causes serve model degradation from both identity similarity and image quality. Across all perturbations, ASPL causes the most severe degradation under the setting without purification, while MetaCloak performs more robustly under rule-based purification. Compared to rule-based purification, diffusion-based approaches achieve better performance in improving both identity similarity and image quality in most settings. Among them, GrIDPure yields relatively better purification performance since it considers the structure consistency, which suppresses the generative nature during the purification. However, there are still gaps in the IMS score compared to the clean case, and most of the Q scores after conducting GrIDPure purification are still negative. Compared to these baselines, our method closes the gap by further improving the IMS and Q scores, which are even higher than the clean training case for most of the settings. The reasons are two folds: first, we use image-restoration-based approaches, which well perverse the structure; and furthermore, our negative token learning and quality-enhanced sampling contribute to the quality improvement.
Perturbed LatentDiffPure DDSPure GrIDPure IMPRESS Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Table 2: Faithfulness and time cost of different diffusion-based purifications.
@@
Efficiency and Faithfulness of Purification. We present the evaluation of time cost and faithfulness compared to all diffusion-based purification in Table 2. The time cost is measured in seconds per sample with consideration of model loading. Compared to other methods, our purification has the lowest time cost and is 10$\times$ faster than the previous SoTA, IMPRESS. The reason behind this is that we leverage the super-resolution module, which empowers the usage of skip-step sampling to boost the generation time. Moreover, we test the purification faithfulness of each method in terms of LPIPS loss [33], a common metric measuring the visual perception distance of two images. From Table 2, we can see that our method achieves the lowest LPIPS loss. To visually validate this, we additionally present the purified images in Figure 5. From the figure, we can see that other diffusion-based approaches have limitations in hallucinating the content, introducing severe artifacts, or not having enough purification strength. In particular, we observed that LatentDiffPure causes a great change in identity during the purification while GrIDPure brings some artifacts. In comparison, our purification greatly improves faithfulness with off-the-shelf image restoration models that are prone to have nearly the same structure input and output.
@@ 5.3 Resilience Against Adaptive Perturbations
Table 3: Different variants under Adaptive Attack (AA).
Method Neg. Token Before AA After AA
IMS Q Avg. IMS Q Avg. $\Delta$
Ours-CodeSR x 0.18 0.41 0.29 -0.06 0.03 -0.01 $\downarrow$0.30
$\checkmark$ 0.23 -0.03 0.10 -0.09 0.24 0.08 $\downarrow$0.02
Ours-Code x 0.14 0.42 0.28 -0.09 -0.17 -0.13 $\downarrow$0.41
$\checkmark$ 0.05 0.11 0.08 -0.16 -0.47 -0.32 $\downarrow$0.40
Ours-SR x 0.07 -0.02 0.03 0.20 -0.39 -0.09 $\downarrow$0.12
$\checkmark$ 0.07 -0.02 0.03 -0.03 0.08 0.02 $\downarrow$0.01
DNN-based purification is prone to further adaptive attacks due to the non-smoothness in terms of latent representation space [11] and also the vulnerability by exploiting adversarial examples [9]. To validate whether our framework can still work upon further adversarial perturbation with new knowledge of our pipeline, we additionally conduct experiments on evaluations of different variants of our approach before and after the adaptive perturbation crafted against the image purification part. The adversarial perturbation is crafted following AdvDM [7] with consideration of the CFG [39] sampling trajectory with a large perturbation budget of $r=16/255$. For the model variants, we consider the full variant with both modules turned on, and also the ablated versions with one of them turned off. From Table 3, we can see that the full variant with Neg. Token is the most effective and robust one under adaptive attack across other variants. Furthermore, we notice that the variant with only SR is also robust to the adaptive attack since it yields the lowest change in average score after the attack. However, only using SR is not enough, since when we remove the CodeFormer, the purification is sub-optimal. These results conclude that our method is robust to further adaptive perturbation and potentially a once-for-all defense approach.
5.4 Ablation Study
Contribution of Modules. We present full ablations on the three modules in our method in Table 4. From the table, our method works best under the full setting. When turning off any of the modules, the average performance degrades, with turning off Neg. Token suffers the most. Surprisingly, when only enabling the SR module, the IMS score is relatively good but with bad quality. For the settings that only enable one of the modules, we find that enabling Neg. Token also yields the best performance. These indicate that Neg. Token is quite important when training on purified images. Furthermore, for the settings that only allow two modules enabled, we found that the SR+Neg.Token yields the best performance, with just some gap between the IMS scores. In conclusion, the results suggest that all the modules contribute to the learning performance gain in IMS and quality.
Table 4: Ablation study.
@@ 5.4 Ablation Study
0.266 0.374 0.320
0.199 0.433 0.316
0.153 0.209 0.181
0.207 0.105 0.156
0.147 0.039 0.093
0.184 -0.042 0.071
0.234 -0.378 -0.072
0.096 -0.634 -0.269
Figure 6: Evaluation of generation performance of models across different stable diffusion versions.
Different Models and Radii. We additionally validate the effectiveness of our method under different LDM versions. From Figure 6, we can see that across different stable diffusion versions, models trained on our purified data yield better performance in terms of IMS, Q, and average scores under most of the settings. And the major improvement is in the quality part. We also demonstrate the effectiveness of our method across different perturbation radii at App. B. These results conclude the generalizability of our method across settings.
6 Conclusion, Limitations, and Future Works
@@
In this paper, we dive into the underlying mechanism behind the effectiveness of existing error-maximizing protective perturbation approaches against the unauthorized fine-tuning of personalized diffusion models. Motivated by the latent mismatch observation, we propose to use super-resolution and image restoration models for latent realignment. Furthermore, we propose contrastive learning with negative tokens based on the analysis from the shortcut learning perspective. Extensive experiments demonstrate the effectiveness, efficiency, and faithfulness of our method.
Despite being mainly evaluated on the facial dataset, our framework is generally designed and can be adapted for various image domains, like objects and artwork. This indicates that it might be used to further invade those already “protected” data with previous perturbations, which might create some negative societal impact but also promote the advancement of this domain. Furthermore, in Section 5.3, we find that CodeFormer might be more vulnerable under adaptive attacks, but it does better in purification than the SR. Thus, there might still be room to improve in determining the optimal layer combination of the SR and CodeFomrer modules to maximize the robustness-effectiveness trade-off.
Appendix A Experiment Details
A.1 Hardware and Training Details
Hardware Details. All the experiments are conducted on an Ubuntu 20.04.6 LTS (focal) environment with 503GB RAM, 10 GPUs (NVIDIA® RTX® A5000 24GB), and 64 CPU cores (Intel® Xeon® Silver 4314 CPU @ 2.40GHz). Python 3.9.18 and Pytorch 1.13.1 are used for all the implementations.
Training and Inference Settings. The Stable Diffusion (SD) v2-1-base [50] is used as the model backbone. For Dreambooth training, we conduct full fine-tuning, which includes both the text-encoder and U-Net model with a constant learning rate of $5\times 10^{-7}$ and batch size of 2 for 1000 iterations in mixed-precision training mode. We use the 8-bit Adam optimizer [51] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ under bfloat16-mixed precision and enable the xformers for memory-efficient training. For calculating prior loss, we use 200 images generated from Stable Diffusion v2-1-base with the class prompt “a photo of a [class norn]”. The weight for prior loss is set to 1. For the evaluation phase, we set the inferring steps as 100 with prompts “a photo of sks person” and “a smiling photo of sks person” during inference to generate 16 images per prompt. For all the settings, the Classifier-free Guidance (CFG) is turned on by default with a guidance scale of $7.5$.
A.2 Metrics
@@
In this section, we describe the evaluation metrics used in our experiments in more detail. Following [8], we use CLIP-IQAC, which calculates the CLIP score difference between “a good photo of [class]” and “a bad photo of [class]”. For calculating IMS-VGGNet, we leverage the VGGNet in the DeepFace library for face recognition and face embedding extraction [52]. For IMS-IP, we leverage antelopev2 model from InsightFace library [43] following IP-adapter [44]. We report the weighted average of them with a weighting factor on IMS-IP as 70% since we find it yields a more stable evaluation with IMS-VGG as 30%. We compute all the mean scores over all generated images and all instances. For the instance $i$ and its $j$-th metric, its $k$-th observation value is defined as $m_{i,j,k}$. For the $j$-th metric, the mean value is obtained with $\sum_{i,k}m_{i,j,k}/(N_{i}N_{k})$, where $N_{i}$ is the instance number for that particular dataset, and $N_{k}$ is the image generation number.
Algorithm 1 Training clean LDM on adversarial perturbed data with systematic defending (Summary of Sec 4.2)
1:Corrupted training set $X^{\prime}_{0}$, LDM $\theta_{0}$, CodeFormer $\phi=\{\mathcal{E}_{\phi},\mathcal{D}_{\phi},\mathcal{T}_{\phi},\mathcal{C}\}$, SR model $\psi$, Prior data $\bar{X}_{0}$, Negative Token $\mathcal{V*}_{N}$, Instance prompt $c^{ID}$, Class prompt $c$, #Generation $N_{gen}$, Quality negative prompt $\bar{\mathcal{P}_{Q}}$, Personalized identifier $\mathcal{V*}$
2:Personalized diffusion model with clean-level generation performances $\theta_{T}$, and clean personalized generation $X_{gen}$
3:// Step 1: Input Purification
4:CodeFormer: Low-quality latent $Z_{l}=\mathcal{E}_{\phi}(X^{\prime}_{0})$Rightarrow Code prediction $\tilde{Z}_{c}=\mathcal{T}(Z_{l},\mathcal{C};\phi)$Rightarrow High-quality restoration $\tilde{X}_{0}=\mathcal{D}_{\phi}(\tilde{Z}_{c})$
5:Super-resolution: Resize $\tilde{X}_{0}$ to $128^{2}$Rightarrow Applying SR model $\psi$ with quality-guided prompt and obtain $\tilde{X}_{0}^{\text{purified}}$ in $512^{2}$
6:// Step 2: Training Personalized LDM with Contrastive Decouple Learning
7:for $i$ in $1,\cdots,T$ training steps do
8:     Sample an instance data $x_{i}$ from $\tilde{X}_{0}^{\text{purified}}$, and an class-prior data $\bar{x}_{0}$ from $\bar{X}_{0}$
9:     Crafting decoupled instance prompt $c^{ID}_{dec}=\texttt{concat}(c^{ID},\mathcal{V*}_{N})$ and class-prior prompt $c_{dec}=\texttt{concat}(\mathbf{c},\text{``without''},\mathcal{V*}_{N})$.
10:     Optimize the LDM $\theta_{i}$ with standard DreamBooth loss $\mathcal{L}_{db}$triangleright Following Eq. 2
11:     $\mathcal{L}_{db}(x_{i},c^{ID}_{dec},\bar{x}_{0},c_{dec};\theta_{i})=\mathcal{L}_{\mathrm{denoise}}\left(x_{i},c^{ID}_{dec}\right)+\lambda\mathcal{L}_{\mathrm{denoise}}\left(\bar{x}_{0},c_{dec}\right),$
12:     Update LDM weight $\theta_{i}$ with $\nabla_{\theta_{i}}\mathcal{L}_{db}$ using AdamW optimizer on UNet Denoiser and Text Encoder
13:end for
14:// Step 3: Noise-free Sampling Process during Inference
15:for $j$ in $1,\cdots,N_{gen}$ do $\triangleright$ Following Eq. 6 with Classifier-free Guidance
16:     Sampling $X_{gen}^{j}$ with personalized identifier $\mathcal{V*}$, and negative prompt of $\text{concat}(\mathcal{V*}_{N},\bar{\mathcal{P}_{Q}})$ with personalized LDM $\theta_{T}$
17:end for
18:return $\theta_{T}$, $X_{gen}$
@@ Appendix B More Experimental Results
Different Radii. We present the evaluation of the models trained on perturbed data and purified data under different perturbation radii in Table 8. As the results shown in the Table, we found that our method consistently yields better models in terms of IMS and Q score. This demonstrates that our method is insensitive to different perturbation radii.

### v4

@@ 5.2 Effectiveness, Efficiency, and Faithfulness
Effectiveness Comparison. We present the effectiveness of different purification across seven perturbation methods in Tab. 1. From the table, we can see that compared to the clean case, training on perturbing data causes serve model degradation from both identity similarity and image quality. Across all perturbations, ASPL causes the most severe degradation under the setting without purification, while MetaCloak performs more robustly under rule-based purification. Compared to rule-based purification, diffusion-based approaches achieve better performance in improving both identity similarity and image quality in most settings. Among them, GrIDPure yields relatively better purification performance since it considers the structure consistency, which suppresses the generative nature during the purification. However, there are still gaps in the IMS score compared to the clean case, and most of the quality scores after conducting GrIDPure purification are still negative. Compared to these baselines, our method closes the gap by further improving the IMS and quality scores, which are even higher than the clean training case in all the settings. The reasons are twofold: first, we use image-restoration-based approaches, which preserve the image structure well; furthermore, our CDL module contributes significantly to quality improvement. Please refer to the App. B for the full comparison results with standard deviations.
Perturbed LatentDiffPure-$\emptyset$ LatentDiffPure DDSPure GrIDPure IMPRESS Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Efficiency and Faithfulness of Purification. We present the evaluation of time cost and purification faithfulness compared to all other diffusion-based purification approaches in Tab. 3. The time cost is measured in seconds per sample with consideration of model loading. Compared to other methods, our purification has the lowest time cost and is 10$\times$ faster than the previous SoTA method, IMPRESS. The reason behind this is that we leverage the super-resolution module, which empowers the usage of skip-step sampling to boost the generation time. Moreover, we test the purification faithfulness of each method in terms of LPIPS loss (Zhang et al., 2018), a common metric measuring the visual perception distance of two images. From Tab. 3, we can see that our method achieves the lowest LPIPS loss. To visually validate this, we additionally present the purified images in Fig. 4. From the figure, we can see that other diffusion-based approaches have limitations in hallucinating the content, introducing severe artifacts, or not having enough purification strength. In particular, we observed that LatentDiffPure causes a great change in identity during the purification, which might be attributed to the semantic distortion during the purification process in latent space. On the other hand, GrIDPure (Zhao et al., 2024) brings some artifacts to the purified image, which indicates that the underlying unconditional diffusion model pre-trained on ImageNet might not be suitable for general domain purification. In comparison, our purification method significantly enhances faithfulness by leveraging off-the-shelf image restoration models. These models are designed to preserve the structural integrity of the input, resulting in output images that closely maintain the original composition while effectively removing perturbations. This approach ensures that the purified images retain the essential features and identity of the original subjects, while successfully mitigating unwanted artifacts or noise.
Table 2: Faithfulness and efficiency of different diffusion-based purifications.
@@
Table 3: Effectiveness of different model variants against Adaptive Attacks (AA).
Modules CDL Before AA After AA $\mathbb{E}$[Avg.]
IMS Q Avg. IMS Q Avg.
CodeSR 0.256 0.514 0.385 0.116 -0.070 0.023 0.204
-0.215 0.028 -0.094 -0.313 -0.533 -0.423 -0.259
Code 0.294 0.385 0.339 0.138 -0.104 0.017 0.178
-0.336 0.020 -0.158 -0.382 -0.474 -0.428 -0.293
SR 0.190 0.260 0.225 0.249 -0.182 0.034 0.130
-0.059 -0.439 -0.249 -0.114 -0.616 -0.365 -0.307
@@ 5.3 Resilience Against Adaptive Perturbations
DNN-based purification is prone to further adaptive attacks due to the non-smoothness in terms of latent representation space (Guo et al., 2023) and also the vulnerability by exploiting adversarial examples (Ilyas et al., 2019). To validate whether our framework can still work upon further adversarial perturbation with new knowledge of our pipeline, we additionally conduct experiments on evaluations of different variants of our approach before and after the adaptive perturbation crafted against the image purification part. The adversarial perturbation is crafted following AdvDM with consideration of the CFG (Ho & Salimans, 2022) sampling trajectory with a large perturbation budget of $r=16/255$. For the model variants, we consider the full variant with both modules turned on, as well as the ablated versions with one of them turned off. From Tab. 3, we can see that the full variant with CDL is robust to the adaptive attack across other variants in terms of performance drop. Furthermore, we notice that the variant with both SR and CDL yields a slightly better average score than the CodeSR configuration after the attack. This indicates that the CodeFormer module might be more susceptible to the adaptive attack while the SR module is more robust. However, using SR with CDL solely in case cases gives sub-optimal purification results. Our CodeSR configuration with CDL gives a better expected overall performance under mixed perturbation scenarios with $P$(AA)=50
5.4 Ablation Study and Sensitivity Analysis
Contribution of Individual Modules. We present ablations on the three modules in our method in Tab. 4. From the table, our method works best under the full setting. When turning off any of the modules, the average performance degrades, with turning off CDL suffers the most. On the other hand, if we only turn on one of the modules, we find that CDL is still the most important one that retains higher generation performance. Furthermore, if we only do input purification without CDL, the generation quality is not as good as the full setting with CDL. This indicates that CDL is crucial for the performance of our method. Surprisingly, when only enabling the SR module, the IMS score is relatively good but with bad quality. While turning on the CodeFormer module alone, the boost is more on the quality score side. The settings that enable these two modules together yield a higher average score. These indicate that SR and CodeFormer modules are complementary to each other. Furthermore, for the settings that only allow two modules enabled, we found that the combination of CodeFormer and CDL yields the best performance compared to the other two combinations. Furthermore, we visualize the quality-score curve of identifier $\mathcal{V}*$ that shows consistent improvement during training in the App. B.1. In conclusion, the results suggest that all modules contribute to the learning performance gain in both IMS and quality scores.
Table 4: Ablation study on individual modules.
@@ 5.4 Ablation Study and Sensitivity Analysis
0.256 0.514 0.385
-0.215 0.028 -0.094
0.294 0.385 0.339
0.190 0.260 0.225
-0.336 0.020 -0.158
-0.059 -0.439 -0.249
0.160 0.038 0.099
-0.271 -0.425 -0.348
Figure 5: Generations from models trained on: (left) clean data, (middle) perturbed data without defense, and (right) purified data using our defense approach. The results demonstrate that our defense method significantly enhances generation quality, bringing it closer to clean data levels.
Generation Visualization and Sensitivity Test. We further visualize the generation of models trained in three cases, including clean, perturbed, and purified in Fig. 5. The visualization demonstrates that our defense greatly helps retain clean-level generation quality. Additionally, we find that the concept learned associated with $\mathcal{V}*$ under perturbed case matches the noise concept learned using CDL alone, indicating the CDL successfully decouples the learning of noise patterns (refer to the App. B.2). The sensitivity analysis of noise tokens is provided in App. B.3.
6 Conclusion
@@
In this paper, we dive into the underlying mechanism behind the effectiveness of existing protective perturbation approaches against the unauthorized fine-tuning of personalized diffusion models. Motivated by the latent mismatch observation, we propose to use super-resolution and image restoration models for latent realignment. Furthermore, we propose contrastive decoupling learning with quality-enhanced sampling based on the analysis from the shortcut learning perspective. Extensive experiments demonstrate the effectiveness, efficiency, and faithfulness of our method. Despite being mainly tested on facial data, our framework can generalize to other domains beyond the facial domain. Future work could optimize module combinations for balanced utility and robustness (in Sec. 5.3), and develop stronger protection methods based on our framework’s robustness-effectiveness trade-off.
Appendix A Implementation Details
A.1 Metrics
@@
In this section, we describe the evaluation metrics used in our experiments in more detail. Following (Liu et al., 2024), we use CLIP-IQAC, which calculates the CLIP score difference between “a good photo of [class]” and “a bad photo of [class]”. For calculating IMS-VGGNet, we leverage the VGGNet in the DeepFace library for face recognition and face embedding extraction (Serengil & Ozpinar, 2021). For IMS-IP, we leverage antelopev2 model from InsightFace library (Deng et al., 2020) following IP-adapter (Ye et al., 2023). We report the weighted average of them with a weighting factor on IMS-IP as 70 We compute all the mean scores for all generated images and instances. For the instance $i$ and its $j$-th metric, its $k$-th observation value is defined as $m_{i,j,k}$. For the $j$-th metric, the mean value is obtained with $\sum_{i,k}m_{i,j,k}/(N_{i}N_{k})$, where $N_{i}$ is the instance number for that particular dataset, and $N_{k}$ is the image generation number.
A.2 Hardware and Training Details
Hardware Details. All the experiments are conducted on an Ubuntu 20.04.6 LTS (focal) environment with 503GB RAM, 10 GPUs (NVIDIA® RTX® A5000 24GB), and 64 CPU cores (Intel® Xeon® Silver 4314 CPU @ 2.40GHz). Python 3.9.18 and Pytorch 1.13.1 are used for all the implementations. Please refer to the supplementary material for the code and environment setup.
Training and Inference Settings. The Stable Diffusion (SD) v2-1-base (Rombach et al., 2022) is used as the model backbone. For Dreambooth training, we conduct full fine-tuning, which includes both the text-encoder and U-Net model with a constant learning rate of $5\times 10^{-7}$ and batch size of 2 for 1000 iterations in mixed-precision training mode. We use the 8-bit Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ under bfloat16-mixed precision and enable the xformers for memory-efficient training. For calculating prior loss, we use 200 images generated from Stable Diffusion v2-1-base with the class prompt ‘‘a photo of a [class norn]’’. The weight for prior loss is set to 1. For the evaluation phase, we set the inferring steps as 100 with prompts “a photo of sks person” and “a smiling photo of sks person” during inference to generate 16 images per prompt. For all the settings, the classifier-free guidance Ho & Salimans (2022) is turned on by default with a guidance scale of $7.5$. For the implementation of baseline methods, please refer to App. D.
@@ Appendix B More Experimental Results
Figure 6: LIQE quality score of $\mathcal{V}*$.
B.1 Quality Score Curve During Training
We present the LIQE (Zhang et al., 2023) quality score curve during fine-tuning under different settings, including clean training, vanilla training on perturbed data, training with CDL, and training with CodeSR+CDL in Figure 6. This curve illustrates the evolution of image quality throughout the training process. As evident from the figure, our proposed decoupled learning (CDL) approach significantly enhances the quality compared to the case with perturbations. Moreover, when we combine CDL with input purification (CodeSR + CDL), the model achieves quality performance comparable to clean-level training. These results further validate the effectiveness of our proposed method in defending against adversarial perturbations and maintaining high-quality outputs in PDMs.
B.2 Learned Concepts Visualization
Figure 7: More visualization on the learned personalized and noise concepts from trained models with and without CDL. With concept extraction, we examine that using CDL successfully decouples the original noise pattern spuriously linked to personalized tokens $V*$ to the noise tokens $\mathcal{V}_{N}^{*}$.
To visually demonstrate our method’s effectiveness, Fig. 8 compares the concept extraction results from trained models with vanilla training, CDL, and CodeSR+CDL. We extract three concepts from the trained models, including the instance concept, instance+class concept, and decoupled noise concept. The third one aims to visualize the noise pattern from the perturbed data that we seek to decouple. From the figure, we can see that CDL helps the model learn the correct concept-image correlations while adding CodeSR, which further improves the generation quality. Interestingly, we find that the learned noise concept in CDL-based training matches the pattern of the one falsely linked by the personalized concept in vanilla training. We present more results supporting this in Fig. 7. This validates the effectiveness of our method in learning the correct concept-image correlations and decoupling the noise concept. Furthermore, from Fig. 8, we find that adding input purification (CodeSR) greatly boosts generation quality. Under the purification case, the contribution of CDL is more about decoupling the left-over background artifacts from the personalized concept.
Vanilla Training (No defense) CDL CodeSR + CDL
Inference the learned Instance Concept $\Rightarrow$ Prompt: “a photo of $\mathcal{V}*$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Inference the learned Instance+Class Concept $\Rightarrow$ Prompt: “a photo of $\mathcal{V}*$ Person
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Inference the Decoupled Noise Concept $\Rightarrow$ Prompt: “a photo of $\mathcal{V}_{N}*$ noisy perturbation
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Concept extraction with three different prompts from the trained model with vanilla training, CDL, and CodeSR+CDL. Results show that CDL alone helps models learn the right correlations, and adding input purification further boosts the generation quality.
B.3 Contrastive Decoupling Learning with Different Noise Tokens.